OpenStudy (anonymous):

Fred lives in Syler City, where the tax rate is 3.5%. His property is assessed at $93,000. What is his tax?

OpenStudy (anonymous):

3.5/100 * 93,000 = 3.5 *930 = 3255

OpenStudy (anonymous):

is there a decimal anywhere

OpenStudy (anonymous):

Nope